In order for applicants to work for the foreign-service department, they must take a test in the language of the country where they plan to work. The data below shows the relationship between the number of years that applicants have studied a particular language and the grades they received on the proficiency exam.

# of years,x 2 3 3 4 5 5 6 6 8
Grade on test,y 59 62 70 72 78 85 80 88 92

Find the equation of the regression line for the given data.

Group of answer choices

= 49.61x + 5.7

= 49.61x - 7.08

= 7.08x - 49.61

= 5.7x + 49.61
Answer #1

ΣX = 42
ΣY = 686
ΣX * Y = 3361
ΣX2 = 224
ΣY2 = 53326

Equation of regression line is Ŷ = a + bX
b = (nΣΧΥ – ΣΧΣΥ)/(ηΣΧ? - (ΣΧ)2)
b = ( 9 * 3361 - 42 * 686 ) / ( 9 * 224 - ( 42 )^{2})
b = 5.702
a =( Σ Y - ( b * Σ X) ) / n
a =( 686 - ( 5.7024 * 42 ) ) / 9
a = 49.611
Equation of regression line becomes Ŷ = 49.61 + 5.7 X

= 5.7x + 49.61